Find the slope and the y-intercept
y=3x-2
The slope is ?
The y-intercept is (0,?)
5y + 2x + 8= 16+2x
Slope is?
y-intercept is 0,?

Answer by rfer(16322)

a)
slope=3
y int=(0,-2)
b)
5y+2x+8=16+2x
5y=8
y=8/5=1.6
slope=0
y int=(0,1.6)